A trumpet emits a sound frequency equivalent to 100Hz. If the person using


the trumpet is moving towards the listener, what would be the perceived


frequency of the listener if the velocity of the sound is 350 m/s and the man


using the trumpet is moving at the rate of 5 m/s?

1
Expert's answer
2021-11-19T10:50:40-0500

We can find the frequency observed by the listener from the Doppler Shift formula:


"f_o=f_s(\dfrac{v}{v-v_s}),""f_o=100\ Hz\times(\dfrac{350\ \dfrac{m}{s}}{350\ \dfrac{m}{s}-5\ \dfrac{m}{s}})=101.45\ Hz."